### v2

# EEGDM: EEG Representation Learning via Generative Diffusion Model

Abstract
While electroencephalogram (EEG) has been a crucial tool for monitoring the brain and diagnosing neurological disorders (e.g., epilepsy), learning meaningful representations from raw EEG signals remains challenging due to limited annotations and high signal variability. Recently, EEG foundation models (FMs) have shown promising potential by adopting transformer architectures and self-supervised pre-training methods from large language models (e.g., masked prediction) to learn representations from diverse EEG data, followed by fine-tuning on specific EEG tasks. Nonetheless, these large models often incurred high computational costs during both training and inference, with only marginal performance improvements as the model size increases. In this work, we proposed an EEG representation learning framework building upon Generative Diffusion Model (EEGDM). Specifically, we developed a structured state-space model for diffusion pretraining (SSMDP) to better capture the temporal dynamics of EEG signals and trained the model using a Denoising Diffusion Probabilistic Model. Subsequently, the resulting latent EEG representations were then used for downstream classification tasks via our proposed latent fusion transformer (LFT). To evaluate our method, we used multi-event datasets covering both interictal epileptiform discharges and seizure detection, and compared EEGDM with current state-of-the-art approaches, including EEG FMs. Empirical results showed that our method outperformed the existing methods. These findings suggested that EEGDM offered a promising alternative to current FMs. Our code is available at: https://github.com/jhpuah/EEGDM.
Index Terms:
Electroencephalogram, diffusion model, representation learning, transformer, structured state-space model.
I Introduction
Electroencephalography (EEG) provides a noninvasive technique for monitoring the brain by recording electrical signals arising from neural activity and pathological processes, with high temporal resolution. Its effectiveness has led to widespread use in brain–computer interfaces (BCIs), neuroscience research, and neurological diagnostics [1, 2]. A key clinical application is the management and treatment of epilepsy, which affects approximately 50 million patients worldwide [3], where EEG recognition plays a critical role in diagnosis and pre-surgical planning. Numerous machine learning and deep learning models have been developed to detect interictal epileptiform discharges (IEDs) and seizures from EEG recordings, which often include various other brain activities (e.g., eye movements, artifacts, and other background activity). Despite showing initial promise, improving classification performance remains an active area of research for clinical adoption.
A major challenge behind the performance bottleneck in IED and seizure detection is the difficulty of learning robust and meaningful representations of EEG signals, due to limited high-quality annotation and high signal variability across subjects, conditions, and brain activities [2, 4]. Recently, inspired by the success of large language models (LLMs), various EEG foundation models (FMs) [4, 5] have been proposed and achieved different degrees of success. Following the procedure of LLMs, EEG FMs were pretrained on a large amount of unlabeled and diverse EEG data through self-supervised pre-training for representation learning using transformer architectures [6]. In most EEG FMs, the signals were tokenized into sequential tokens, with a subset masked to enable self-supervised training via masked token prediction [1, 2, 7]. These methods encouraged the models to learn the underlying relationships (through self- and cross-attention modules in the transformer) between tokens and effectively capture meaningful representations without relying on annotation. Subsequently, these FMs were fine-tuned on specific downstream tasks such as IED detection [2], sleep stage classification [8], and emotion recognition [9]. Although they often outperformed traditional deep learning models, they incurred high computational costs during both training and inference, with only marginal performance improvements as model size increased.
Diffusion models [10] represent another class of self-supervised learning and generative models that have demonstrated remarkable success in image [10, 11] and audio generation [12, 13]. Recent studies suggested that diffusion models also hold strong potential for representation learning tasks [14]. Unlike masking-based methods, diffusion models achieve self-supervision by progressively corrupting input data with Gaussian noise in a forward process and learning to recover the original data through a reverse denoising process. Several approaches [15, 16, 17] have been proposed to leverage diffusion models to extract representations for classification and segmentation of visual tasks. However, the generalization and applicability of diffusion models in EEG representation learning remain underexplored, as most existing works primarily employed them for signal generation and data augmentation [18, 19, 20], rather than for representation learning. Furthermore, it is also unclear whether the learned representation would be useful and how to use it for downstream classification tasks. Hence, this work attempts to fill this gap.
In this work, we propose an EEG Representation Learning Framework, building upon the Generative Diffusion Model (EEGDM). Specifically, we developed a neural architecture based on a structured state-space model for diffusion pretraining (SSMDP) to better capture the temporal dynamics of EEG signals and trained the architecture using a Denoising Diffusion Probabilistic Model (DDPM). Subsequently, the resulting latent EEG representations were then used for downstream classification tasks via our proposed latent fusion transformer (LFT). To evaluate our method, we used the Temple University EEG Event Corpus (TUEV) [21] and Children’s Hospital Boston-Massachusetts Institute of Technology (CHB-MIT) [22], which contained IED and seizure, respectively, and compared EEGDM with current state-of-the-art approaches, including EEG FMs.
The main contributions of this work are as follows:
  • We presented EEGDM, a diffusion model-based framework for learning EEG signal representations and classification of multi-event EEG, extending diffusion model beyond signal generation and data augmentation.
  • We developed SSMDP to capture the temporal dynamics of EEG signals and trained it via the forward and reverse process of DDPM for representation learning using TUEV dataset.
  • We proposed LFT to leverage and integrate the latent representations learned by SSMDP for downstream tasks, including TUEV (seen during pretraining) and CHB-MIT (unseen during pretraining) datasets.
  • We empirically compared our method with state-of-the-art on TUEV and CHB-MIT, and provided a detailed ablation study to analyze its components.
II Related Works
This section reviews the evolution of EEG representation learning, from classical supervised, unsupervised, and semisupervised approaches to recent advances in self-supervised learning (particularly FMs and diffusion models) and their emerging potential for EEG, with an emphasis on the latter.
II-A Supervised and Unsupervised Approaches
EEG representation learning has often arisen as a byproduct of supervised approaches, such as the seminal common spatial pattern methods [23] and more recent deep neural network models [24]. Such representations have been examined across various neural architectures, including convolutional neural networks [25], graph neural networks [26, 27], and Transformers [28, 29], revealing findings that aligned with known brain activity patterns. When annotations are unavailable, unsupervised approaches, particularly autoencoders and their variants, offered an alternative. Variational autoencoder (VAE) has shown effectiveness in learning robust representation in the regularized latent space, achieving transfer learning from motor execution signal to motor imaginary signals [30]. Another approach was metric learning, which embedded EEG signals into a vector space where similarity was defined as proximity between samples, typically optimized using contrastive loss (e.g., [31]). However, these methods are constrained by the availability of annotation and often yield task-specific representations that lack generalizability.
Another line of research was semi-supervised learning that leverages a small amount of annotated data to train a model, while using the model itself to generate “pseudo-labels” for unannotated data, which can then be used as additional training data. For example, in [32], the empirical probabilities of the training data are used as pseudo-labels to calibrate model certainty and confidence, thereby addressing the scarcity of annotations in motor imagery classification. Another promising approach was self-supervised learning, which learns by corrupting and subsequently restoring the training data. Common corruption strategies included masking, adding noise, or combining both. These techniques have been successfully applied to various EEG tasks, such as sleep stage classification [33], evoked response enhancement [34], and motion intention recognition [35]. Recently, self-supervised pretraining methods, particularly mask prediction with transformers and diffusion models, have demonstrated strong potential across multiple domains. However, only a few pioneering studies have begun to explore their effectiveness in EEG representation learning.
II-B EEG FMs
Inspired by LLMs, EEG FMs employed token-based and masking-based self-supervised pretraining on a diverse and large amount of unannotated data with transformer architectures before finetuning the pretrained model with annotation in a supervised manner. The pre-training allowed EEG FMs to grasp the underlying structure of EEG from various problem domains and experimental paradigms, thus addressing the annotation scarcity issue and reducing the risk of overfitting [36]. Prior to pre-training, the raw EEG signal was usually preprocessed and then segmented into patches. A tokenizer, usually a variant of vector-quantized (VQ) VAE [37], was trained to encode these patches into discrete representations, known as tokens. The VQ encoder first projected a patch to a vector representation, then the representation was replaced by the most similar entry (token) from a codebook. The VQ decoder then reconstructed a target from the token, either the original patch itself (in CBraMod [6] and Gram [7]) or a transformation of the original patch, such as Fourier spectrum (in LaBraM [2]). This process ensured that the tokens in the codebook capture semantically meaningful information about the original patches.
Figure 1: Illustration of the proposed SSMDP for representation learning. The model learned to denoise and reverse a noised single-channel EEG signal by predicting the diffusion velocity, conditioned on diffusion step embedding and channel embedding, using S4D architecture. The latent activities from the gate or filter channels were then leveraged for downstream classification.
Subsequently, self-supervised pretraining operates on the sequence tokenized by the VQ encoder and the codebook. Initially, a subset of a sequence is masked by randomly replacing some of the tokens with a “mask token”. The masked sequence is then processed by a transformer-based model to predict the index of the replaced tokens in the codebook. The neural architecture of the model varies: LaBraM [2] used a classic transformer, CBraMod [6] employed a criss-cross attention mechanism to model spatial and temporal dependencies in parallel, and Gram [7] used a masked autoencoder consisting of a layer-fusion encoder and a transformer-based decoder. Subsequently, the pretrained model can be finetuned for classification tasks. Usually, a trainable “class token” is concatenated into the input sequence to be processed by transformer. Then, a linear classification head was trained to classify the processed class token to a label. While EEG FMs often outperformed classical deep learning models, they incurred high computational costs during both training and inference, with only marginal performance improvements as model size increases.
II-C Diffusion Model in Representation Learning
Diffusion models are closely related to denoising autoencoders [38], which are capable of producing meaningful representations in their latent space. In particular, denoising diffusion autoencoder [14] showed that the inner working of diffusion models was analogous to an autoencoder. Several works explored the efficiency of latent activities of diffusion models in image classification [16, 17] and image segmentation task[15].
In addition to latent activities, the connection between conditional diffusion models and conditional probability distributions also led to alternative ways in representation learning. Diffusion classifier [39] showed that any conditional diffusion model can be converted into a classifier without additional training. Joint diffusion model [40] unified this generator-classifier duality: the model was explicitly trained to minimize both classification loss and generative loss, so it can perform both tasks in a single model. These findings presented a promising direction for generalization to EEG representation learning and classification tasks, despite being underexplored.
II-D Diffusion Models for EEG Signals
There is limited work on adapting diffusion models to EEG signal representation learning. Early works have primarily used EEG signals as conditioning inputs to generate other modalities, such as images [41] and audio [42], rather than learning representations of the EEG itself. A more relevant line of research [43, 44] integrates diffusion models within an autoencoder framework, so that the reconstruction errors from the diffusion model can help formulate a more informative latent space, thereby enhancing performance in imagined speech classification [43] and driving behavior classification [44]. Inspired by image super-resolution (e.g. [45]), [46] proposed a diffusion model that can generate high-resolution EEG from low-resolution EEG, to migrate the data scarcity issue for downstream classification task and improve spatial resolution for signal source localization. [47] applied a conditional diffusion model to generate a clean EEG signal under the guidance of noisy EEG signals, hence suppressing EEG artifacts such as electrooculography artifacts.
Several works [18, 19, 20] quantitatively evaluated the similarity between the generated and real signals, and used the synthesized data for augmentation in various downstream tasks, such as imagined motor activity detection and major depressive disorder classification. Nonetheless, the generalization and applicability of diffusion models in EEG representation learning remain underexplored. Moreover, the optimal strategies for leveraging the learned representations are still unclear. These open questions motivate the present work.
Figure 2: From the pretrained SSMDP, we obtain the latent activities of each EEG channel as a four-dimensional tensor: (# of channels, # of layers, # of time points, # of gate channels per layer). The latent activities are then pooled along the time dimension, reducing the dimensionality to (# of channels, # of layers, # of pools, # of gate channels per layer). Colors indicate the latent activities from different channels.
III Methodology
We first present the preliminaries of key components (diffusion models, state-space models, and transformers) that EEGDM builds upon, followed by our EEGDM.
III-A Preliminaries
III-A1 Diffusion Model
A diffusion model consists of a forward process and a reverse process. The forward process gradually transforms a target data distribution into a standard Gaussian distribution, while the reverse process learns to denoise step by step to reconstruct the original data. After training, the reverse process can be used for generation by gradually transforming a sample from the Gaussian distribution into one that resembles the target data distribution.
The forward process models the trajectory of a data point $\mathbf{x}_{0}$ drifts and diffuses from its original distribution $p(\mathbf{x}_{0})$ to the Gaussian distribution $\mathcal{N}(0,I)$. This process is split into $T$ diffusion steps, where each step is an intermediate distribution dependent on the original data and a noise schedule $\bar{\alpha}$:
$p(\mathbf{x}_{t}|\mathbf{x}_{0})=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ (1)
where $t=1,2,3\dots,T$, $\epsilon\sim\mathcal{N}(0,I)$, and $\bar{\alpha}_{t}\in(0,1)$ is a set of strictly decreasing hyperparameters controlling the noise schedule.
The reverse process starts from a pure Gaussian distribution $q(\mathbf{x}_{T})$. The model is trained to perform denoising under the assumption that each transition step in the reverse process is also Gaussian. This assumption becomes increasingly valid as the number of diffusion steps $T$ grows large. The reverse process is modeled as:
$q_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}(\mathbf{x}_{t-1};\mu(\mathbf{x}_{t},t),\sigma_{t}^{2})$ (2)